OpenStudy (anonymous):

Find the measure of angle X to the nearest degree. (Hint: Use the Law of Sines.) 1) 40 degrees 2) 35 degrees 3) 26 degrees 4) 32 degrees

OpenStudy (anonymous):

Is there a picture?

OpenStudy (mrnood):

it gives you a big hint. Do you know (and understand) the Law of Sines ?

OpenStudy (anonymous):

not really .-. I'm kind of lost

OpenStudy (mrnood):

 \[\frac{ a }{ \sin(A) } = \frac{ b }{ \sin(B) } =\frac{ c }{ \sin(C) }\] That is all the info you need to solve the given question